OpenStudy (anonymous):

what is the slope of a line with the points (2,12) and (6,11)?

OpenStudy (anonymous):

are you familiar with the slope formula ?

OpenStudy (anonymous):

yes i answered it i just want to make sure i struggle with math

OpenStudy (nirmalnema):

y2-y1 / x2-x1

OpenStudy (anonymous):

-0.25

OpenStudy (nirmalnema):

11-12 / 6-2 -1/4

OpenStudy (anonymous):

ok thats what i answered thank you

OpenStudy (uri):

\[m=\frac{ (y2-y1) }{(x2-x1) }\] Slope formula

OpenStudy (anonymous):

m = y2 - y1 / x2 - x1 (2,12) x1 = 2 and y1 = 12 (6,11) x2 = 6 and y2 = 11 m = (11- 12) / ( 6 - 2) m = -1/4 yea, thats what I got